Find the equation if the line passing through (-3,2) and having an x-intercept of 1.
Answer by stanbon(75887)

Find the equation if the line passing through (-3,2)
and having an x-intercept of 1.
---------
Form: y = mx+b
If x-intercept is (1,0) 1 = m*0 + b
b = 1
---
So, using (-3,2), solve for "m":
2 = m*-3 + 1
-3m = 1
m = -1/3
------
Equation:
y = (-1/3)x + 1
===================
Cheers,
Stan H.